Please help me solve this.
Examine the rise in gasoline prices from 1997 to 2006. The price of regular unleaded gasoline in January 1997 was $1.26 and in January 2006 the price of refualr unleaded gasoline was $2.31. Use the coordinates (1997,1.26) and (2006. 2.31) to find the slope (or rate change) between the two points. Descruibe how you arrived at your answer.

Answer by NYC Math Tutor(4)

The formula for slope is rise/run or (y2-y1)/(x2-x1)
So, in this case:
(2.31 - 1.26) / (2006 - 1997) = 1.05 / 9 = 0.11666666666
0.11666666 is your slope.